Stone is dropped from the top of the tower and travel 25 m in last second of its motion. If (g=10), height of the tower is

Grade:7

1 Answers

Madhav Verma
6 years ago
Hello Friend.There is shortcut to do this.When a body is in free fall then In 1st sec of dropped of stone it moves 5 mIn 2nd sec of dropped of stone it moves 15 mIn 3rd sec of dropped of stone it moves 25mSo in free fall a body moves in particular sec according to this ratio5:15:25:---------------- an so on1:3:5:7:9:11--------and so onSo according to the question in last sec stone moves 25m It means that body moves 25 m in 3rd sec a/c to questionSo therefore stone moves for 3sec Now solve the question.u=0m/sect=3 seca=g=10m/sec^2S=ut+1/2at^2S=1/2at^2 (u=2m/sec)S=1/2×10×3^2S=45mUse the above shortcut for every free fall object..